# DAL: A Practical Prior-Free Black-Box Framework for Non-Stationary Bandit Environments

### v2

@@ 4.2 On Order-Optimality in Piecewise Stationary Environments
The assumption on the minimum separation between change-points essentially requires scaling as $\tilde{\mathcal{O}}(\sqrt{T/k})$. However, this condition primarily emerges from a conservative proof technique used in [27], where missed detections are aggregated into a single adverse event. Practically, and as corroborated by our experiments in Section 3, this assumption is often violated without negatively impacting the regret performance—even under scenarios with frequent and arbitrarily placed change-points (e.g. $\xi=0.4$). We suspect that this resilience arises because the GLR test, while potentially missing isolated short intervals, reliably detects subsequent changes when stationary segments exceed the threshold length. Even if a change is entirely missed during a segment shorter than $\tilde{\mathcal{O}}(\sqrt{T/k})$, the resulting regret remains under that order. Conversely, when the assumption holds, optimal regret is provably guaranteed. Thus, the required separation threshold acts as a practical ”sweet spot”: segments longer than this threshold are detected reliably, ensuring optimal performance, while shorter segments incur minimal regret, thereby preserving overall optimal regret guarantees.
In line with the black-box design philosophy, Corollary 4.4 enables regret upper bounds across all settings considered, with flexibility in the choice of stationary bandit algorithms. When using specific stationary algorithms from Section 3, DAL matches the state-of-the-art regret bounds in PS-LBs and PS-GLBs at $\tilde{\mathcal{O}}(d\sqrt{N_{T}T})$. More notably, DAL improves the best known bounds in the PS-SCB and PS-KB settings. For PS-SCBs, the strongest, prior-based, bound is due to WeightUCB [57], which achieves $\tilde{\mathcal{O}}(d^{2/3}T^{2/3}N_{T}^{1/3})$ [^4] While MASTER [58] may be extendable to PS-SCBs, no regret bound is currently known.. DAL improves this to $\tilde{\mathcal{O}}(d\sqrt{N_{T}T})$ with our algorithmic choices. Although this matches the bound in [44], their analysis relies on substantially stronger assumptions than those in [27]. For PS-KBs, the prior-free ADA-OPKB [26] achieves $\tilde{\mathcal{O}}(\sqrt{d\gamma_{T}N_{T}T})$, while DAL improves this to $\tilde{\mathcal{O}}(\sqrt{\gamma_{T}N_{T}T})$. This highlights the interesting feature of DAL: the order-wise dependence on problem parameters from the stationary setting seamlessly transfers to the PS setting without degradation. A more detailed comparison of regret bounds is provided in the Appendix.
@@ 4.3 On Drifting Environments
[…]
@@ 5 Summary and Outlook
In this work, we introduced DAL, a practical, prior-free black-box framework for NS bandits. Its plug-and-play design allows seamless integration with a wide range of stationary bandit algorithms. Through extensive experiments in both PS and drifting settings—spanning synthetic and real-world benchmarks—DAL consistently outperforms all prior-free baselines, including the black-box gold standard MASTER and the state-of-the-art ADA-OPKB. DAL even surpasses leading prior-based methods like WeightUCB and WGP-UCB. Finally, DAL achieves the strongest performance in real-world scenarios, which underscores its value as a practical and effective solution.
To explain these results, we provided insights on its theoretical nature. Specifically, using existing theoretical results and providing novel techniques, we showed that DAL achieves order-optimal regret in piecewise stationarity if its input stationary bandit algorithm has order-optimal stationary regret guarantees, under a mild assumption on the separatedness of the change-points. Crucially, DAL preserves and adapts the regret properties of its stationary input algorithm to the PS setting. As a result, it matches the best existing bounds in PS-LBs and PS-GLBs while improving the best known bounds for PS-SCBs and PS-KBs. Regarding drifting non-stationarity, we hypothesized key conditions under which DAL excels–an insight further validated through additional experiments under drifting settings. Our results suggest that a well-designed algorithm for the PS setting can extend to a broad range of drifting scenarios, bridging the gap between these two regimes.
While our work makes significant progress, it also opens several new directions. First, theoretical guarantees for change detection-based methods under drifting environments remain largely unexplored. Second, the current regret bounds for DAL rely on a separation condition between change-points—a standard assumption in the detection-based literature (see e.g., [5, 7, 27])—which nonetheless limits the extent to which DAL achieves fully prior-free theoretical optimality. Addressing these limitations could advance our understanding of detection-based methods in more continuous forms of non-stationarity. Finally, DAL’s modular nature invites future extensions to even richer settings, including general NS reinforcement learning. We believe that deepening the study of piecewise stationarity may be the key to tackling these broader challenges and DAL can serve as a solid foundation towards that goal.
@@ A.1 Change Detectors
The Gaussian GLR test is employed in the case of $\sigma^{2}-$sub-Gaussian distributions, e.g., bounded distributions in $[a,b]$ for some $a<b\in\mathbb{R}$.
@@ A.2 On Forced Exploration in Finite Action Spaces
In practice, the covering set $\mathcal{A}_{\mathrm{e}}$ would generally be selected according to Propositions 4.1 and 4.2, along with the specifications of Corollary 4.4. However, in finite-action settings, the full construction may not be feasible: $\mathcal{A}$ might not contain enough elements to satisfy the required conditions. For instance, in the PS-PB setting, $\mathcal{A}$ may not include $d$ linearly independent actions, while in the PS-KB case, it may lack the full $\delta_{T}$-covering net for the chosen $\delta_{T}$ in Corollary 4.4.
One might expect that when $|\mathcal{A}_{\mathrm{e}}|<d$ in PS-PBs or $|\mathcal{A}_{\mathrm{e}}|<\gamma_{T}$ in PS-KBs, the inability to detect all possible changes would degrade DAL’s performance. In general, however, the algorithm does not need to restart if the changes in the reward function does not change the mean reward obtained from each action. Crucially, as discussed in Appendix B.5, DAL retains order-optimality even in these constrained regimes.
Accordingly, when $|\mathcal{A}|<d$ or $|\mathcal{A}|<\gamma_{T}$, we simply set $\mathcal{A}_{\mathrm{e}}=\mathcal{A}$. In our experiments, the action set is finite (typically in the hundreds). For PS-PBs, the random generation of actions almost always guarantees $d$ linearly independent vectors. For PS-KBs, since $\gamma_{T}$ is typically large, we also use the full action set $\mathcal{A}$ as $\mathcal{A}_{\mathrm{e}}$ without impacting performance.
@@ A.3 Real-World Data Preprocessing
Yahoo! R6A Dataset Construction.
For the NS bandit benchmark based on the Yahoo! R6A click log dataset [^5] Yahoo! Front Page Today Module User Click Log Dataset: https://webscope.sandbox.yahoo.com. we follow the main procedure provided in [10, 62]. We merge ten consecutive days of logs and we group the data by article ID and compute smoothed click-through rates (CTRs) using centered rolling averages over a 100-round window. This generates a time series of empirical CTRs for each article. We segment the dataset into ten distinct subperiods (each spanning half a day), filtering out actions with missing data or high noise. We further select a set of common actions present in all segments to ensure consistent tracking. We average CTRs within each subperiod and smoothing small deviations below a threshold $0.005$. We stack selected actions across multiple days into a single $K\times T$ matrix, where $K$ is the number of valid actions and $T$ the compressed time horizon. To reduce spurious noise and compress the time scale, we apply local smoothing. Finally, we apply post-processing filters to remove (i) globally high-value actions (outliers with inflated CTRs), and (ii) actions that persist as best for too many segments.
Figure 4: Mean rewards for the Yahoo! R6A dataset.
@@
Regarding the stock market experiments we follow the procedure of [17]. For the first experiment, we use the data provided in [17]. For the other experiment, we collect daily closing prices of NASDAQ-100 companies using the Yahoo Finance API [^6] Data retrieved from Yahoo Finance using the publicly available yfinance package. Used solely for non-commercial, academic research purposes.. We filter out stocks with fewer than $T=2000$ trading days and align all time series over the most recent $T$ dates. From this pool, we remove stocks with extremely high volatility or mean price to make the problem non-trivial, then select the top $K$ most volatile stocks from the remainder. In both cases, the stock prices are scaled accordingly to lie in $[0,1]$. Each selected company’s scaled closing price series defines the mean-reward sequence for one arm in a $K$-armed bandit problem. Finally, we corrupt the reward at each time step with $\mathcal{N}(0,0.01)$ noise.
@@
Live Traffic Data Construction.
We construct a NS bandit environment based on the Criteo live traffic dataset [18], following the preprocessing approach of [45] but modeling the problem as an NS-GLB rather than an NS-LB. Specifically, the dataset includes banners shown to users, associated contextual variables, and whether each banner was clicked. We retain the categorical variables cat1 through cat9, along with campaign, which uniquely identifies each campaign. These categorical features are one-hot encoded, and a dimensionality reduction via Singular Value Decomposition selects 50 resulting features. The parameter vector $\theta^{\star}$ is estimated using logistic regression. Rewards are then generated from this regression model with added Gaussian noise of variance $\sigma^{2}=0.01$. Unlike [45], in which the authors employ a single change, we introduce shifts in $\theta^{*}$ via a geometric change-point model with parameter $\xi=0.8$, chaging $60\%$ of the $\theta^{*}$ coordinates at each time-step to $-\theta^{*}$ and extend the horizon to $T=50000$.
@@
We use the Bioliq dataset from [31], comprising a week of readings from 20 power plant sensors. Following their setup, we construct an NS-SCB environment with 190 actions: the reward is $1$ if the last 1000 measurements exceed 2.04, and 00 otherwise. Evaluation is based on cumulative reward. Data available at https://github.com/edouardfouche/G-NS-MAB/tree/master/data.
@@
We evaluate on the Zozo environment, a real-world ad recommender system from [46], using the preprocessed dataset of [31]. We construct an NS-GLB environment with all 80 ads (unlike their 10-action setup), assigning reward $1$ to any ad clicked within one second, and 00 otherwise. Evaluation is based on cumulative reward. Data available at https://github.com/edouardfouche/G-NS-MAB/tree/master/data.

### v4

@@ 4.2 On Order-Optimality in Piecewise Stationary Environments
The assumption on the minimum separation between change-points essentially requires scaling as $\tilde{\mathcal{O}}(\sqrt{T/k})$. However, this condition primarily emerges from a conservative proof technique used in [31], where missed detections are aggregated into a single adverse event. Practically, and as corroborated by our experiments, this assumption is often violated without negatively impacting the regret performance—even under scenarios with frequent and arbitrarily placed change-points (e.g. $\xi=0.4$). We suspect that this resilience arises because the GLR test, while potentially missing isolated short intervals, reliably detects subsequent changes when stationary segments exceed the threshold length. Even if a change is entirely missed during a segment shorter than $\tilde{\mathcal{O}}(\sqrt{T/k})$, the resulting regret remains under that order. Conversely, when the assumption holds, optimal regret is provably guaranteed. Thus, the required separation threshold acts as a practical ”sweet spot”: segments longer than this threshold are detected reliably, ensuring optimal performance, while shorter segments incur minimal regret, thereby preserving overall optimal regret guarantees.
State-of-the-art Regret.
In line with the black-box design philosophy, Corollary 4.5 enables regret upper bounds across all settings considered, with flexibility in the choice of stationary bandit algorithms. When using specific stationary algorithms from Section 3, DAL matches the state-of-the-art regret bounds in PS-LBs and PS-GLBs at $\tilde{\mathcal{O}}(d\sqrt{N_{T}T})$. In PS-CBs, DAL achieves the state-of-the-art regret bound of $\tilde{\mathcal{O}}(\sqrt{|\mathcal{A}|N_{T}T\log|\Pi|})$. More notably, DAL improves the best known bounds in the PS-SCB and PS-KB settings. For PS-SCBs, the strongest, prior-based, bound is due to WeightUCB [61], which achieves $\tilde{\mathcal{O}}(d^{2/3}T^{2/3}N_{T}^{1/3})$. [^5] While MASTER may be extendable to PS-SCBs, no regret bound is currently known. DAL improves this to $\tilde{\mathcal{O}}(d\sqrt{N_{T}T})$ with our algorithmic choices. Although this matches the bound in [49], their analysis relies on substantially stronger assumptions than those in [31]. For PS-KBs, the prior-free ADA-OPKB [30] achieves $\tilde{\mathcal{O}}(\sqrt{d\gamma_{T}N_{T}T})$, while DAL improves this to $\tilde{\mathcal{O}}(\sqrt{\gamma_{T}N_{T}T})$. This highlights the interesting feature of DAL: the order-wise dependence on problem parameters from the stationary setting seamlessly transfers to the PS setting without degradation. A more detailed comparison of regret bounds is provided in the Appendix.
@@ 4.3 On Drifting Environments
$\begin{aligned} &X_{t+1}=\langle\theta_{t},a\rangle+\langle\zeta_{t+1},a\rangle+\varepsilon_{t+1}.\end{aligned}$
@@ 5 Summary and Outlook
We introduced DAL, a practical, prior-free black-box framework for general non-stationary bandits. Its plug-and-play design integrates seamlessly with a wide range of stationary bandit algorithms. Through extensive experiments in both PS and drifting settings—spanning synthetic and real-world benchmarks, DAL consistently outperforms all prior-free baselines, including the black-box gold standard MASTER and the state-of-the-art methods ADA-OPKB and ADA-ILCTB+, and even surpasses leading prior-based methods like WeightUCB and WGP-UCB. Its leading performance in real-world scenarios highlights its value as a practical and effective solution.
On the theoretical side, using existing results and providing novel techniques, we showed that DAL inherits and adapts the regret guarantees of its stationary input algorithm, achieving order-optimal regret under piecewise stationarity, with mild change-point separation. As a result, it matches the best existing bounds in PS-LBs, PS-GLBs and PS-CBs while improving the best known bounds for PS-SCBs and PS-KBs. Regarding drifting non-stationarity, we hypothesized key conditions under which DAL excels–an insight further validated through additional experiments under drifting settings. Our results suggest that a well-designed algorithm for the PS setting can extend to a broad range of drifting scenarios, bridging the gap between these two regimes.
While DAL advances both theory and practice, it opens new directions. First, regret guarantees for detection-based methods in drifting environments remain unexplored. Second, the current regret bounds for DAL rely on a separation condition between change-points—a standard assumption in the detection-based literature (see e.g., [7, 9, 31])—which nonetheless limits the extent to which DAL achieves fully prior-free theoretical optimality. Addressing these gaps would deepen our understanding of detection-based methods in more continuous forms of non-stationarity. Finally, DAL’s modular nature invites extensions to broader settings, including general non-stationary reinforcement learning. We believe that deepening the study of piecewise stationarity may be the key to tackling these broader challenges and DAL can serve as a solid foundation towards that goal.
Acknowledgments
This work was supported in part by a grant from the C3.ai Digital Transformation Institute, by the Army Research Laboratory under Cooperative Agreement W911NF-17-2-0196, and by the US National Science Foundation under grants 2038775 and 2106727.
@@ A.1 Change Detectors
The Gaussian GLR test is employed in the case of $\sigma^{2}$-sub-Gaussian distributions, e.g., bounded distributions in $a,a+2\sigma]$ for some $a\in\mathbb{R}$.
@@ A.2 On Forced Exploration in Finite Action Spaces
Covering Set Construction.
In practice, the covering set $\mathcal{A}_{\mathrm{e}}$ is selected according to Propositions 4.1, 4.2, and Remark 4.3 together with the specifications of Corollary 4.4. However, in finite-action settings, the full construction may not be feasible: the action set $\mathcal{A}$ may not contain enough elements to satisfy the required conditions. For instance, in the NS-PB setting, $\mathcal{A}$ may not include $d$ linearly independent actions, while in the NS-KB case, it may lack a full $\delta_{T}$-covering net for the chosen $\delta_{T}$ in Corollary 4.4. One might expect that when $|\mathcal{A}_{\mathrm{e}}|<d$ in PS-PBs or $|\mathcal{A}_{\mathrm{e}}|<\gamma_{T}$ in PS-KBs, the inability to detect all possible changes would degrade DAL’s performance. In practice, however, DAL does not need to restart when changes in the reward function leave the mean reward of each action unchanged. Crucially, as discussed in Appendix B.5, DAL retains order-optimality even in these constrained regimes. Accordingly, whenever $|\mathcal{A}|<d$ or $|\mathcal{A}|<\gamma_{T}$, we simply set $\mathcal{A}_{\mathrm{e}}=\mathcal{A}$. In our experiments, the action set is finite (typically in the hundreds). For PS-PBs, the random generation of actions almost always guarantees $d$ linearly independent vectors. For PS-KBs, since $\gamma_{T}$ is typically large, we also use the full action set $\mathcal{A}$ as $\mathcal{A}_{\mathrm{e}}$ without impacting performance. On the other hand, since the regret bounds in PS-CBs include $|\mathcal{A}|$, as it is finite, in any PS-CB setting we can simply set $\mathcal{A}_{e}=\mathcal{A}$.
Practical Implementations.
For NS-PBs, we construct $\mathcal{A}_{\mathrm{e}}$ by greedily selecting linearly independent actions until we obtain $d$ such vectors, where $d$ is the dimension of the action space. In the NS-KB setting, $\mathcal{A}_{\mathrm{e}}$ is formed by building a $\delta_{T}$-cover over the bounded action space and choosing the centers of the covering balls. If the action space is continuous and bounded, these centers suffice to cover the space. If the action space is finite and $N_{e}<d^{2p}\gamma_{T}^{2q}\gamma_{T}$, then the entire set $\mathcal{A}$ serves as the covering set, as established in Corollary 4.4. Otherwise, if $N_{e}>d^{2p}\gamma_{T}^{2q}\gamma_{T}$, we select the $d^{2p}\gamma_{T}^{2q}\gamma_{T}$ actions closest to the covering-ball centers. Finally, in the NS-CB setting, selecting a smaller $\mathcal{A}_{e}$ compared to $\mathcal{A}$ does not affect regret, but improves practical performance due to less forced exploration. Thus, depending on the reward function and action set structures, it is recommended to decrease the cardinality of $\mathcal{A}_{e}$ as much as possible.
Experimental Choices.
In our experiments, for NS-PBs the action set is sampled from a multivariate Gaussian distribution, which ensures the existence of $d$ linearly independent actions. Thus, we always set $N_{e}=d$ using the greedy selection procedure described above. For NS-KBs, the regret bound for $N_{e}$ obtained from Theorem 4.3 and Corollary 4.4 is extremely large for our horizons, implying that $|\mathcal{A}|<\gamma_{T}$. Consequently, in all NS-KB experiments we simply take $\mathcal{A}_{\mathrm{e}}=\mathcal{A}$ and set $N_{e}$ equal to the number of available actions, which yielded optimal performance. Finally, since the reward does not exhibit any structure with the arms in PS-CBs, we simply set $\mathcal{A}=\mathcal{A}_{e}$.
@@ A.3 Real-World Data Preprocessing
Microarchitecture Prefetcher Selection Benchmark.
We introduce a non–stationary bandit dataset derived from the MICRO’23 study of [29], built on the SPEC06/17 benchmark suites. Each action corresponds to one of $11$ L2 prefetcher configurations (next–line on/off, stream degree, stride degree). The sequence spans T=26224T{=}26224 rounds; at round $t$, the reward is the trace–level normalized instructions–per–cycle in $[0,1]$, computed from performance counters. We obtained the data directly from the original authors, and note that reproducing the exact series from scratch is not feasible without the same stack, microarchitectural parameters, and arm schedules described in the paper. We aim to release the dataset to facilitate real-world experimentation by the bandit research community.
Figure 4: IPC of the prefetchers of the dataset over time.
@@
Regarding the stock market experiments we follow the procedure of [19]. For the first experiment, we use the data provided in [19]. For the other experiment, we collect daily closing prices of NASDAQ-100 companies using the Yahoo Finance API. [^6] Data retrieved from Yahoo Finance using the publicly available yfinance package. Used solely for non-commercial, academic research purposes. We filter out stocks with fewer than $T=2000$ trading days and align all time series over the most recent $T$ dates. From this pool, we remove stocks with extremely high volatility or mean price to make the problem non-trivial, then select the top $K$ most volatile stocks from the remainder. In both cases, the stock prices are scaled accordingly to lie in $[0,1]$. Each selected company’s scaled closing price series defines the mean-reward sequence for one arm in a $K$-armed bandit problem. Finally, we corrupt the reward at each time step with $\mathcal{N}(0,0.01)$ noise.
@@
COVID–NMA Clinical Dataset Construction.
For the clinical benchmark based on the public COVID-NMA pharmacological RCT database [10], [^7] Data available at: https://doi.org/10.5281/zenodo.14965887. we use only released arm-level counts and metadata and discretize time into calendar months, assigning each trial arm to its Start_date (falling back to Pub_date_online); rows with invalid or missing dates are discarded. We deterministically map case-insensitive rules on treatment type into $13$ actions: Antivirals (any), Anti–inflammatory (steroids/NSAIDs), Interleukin inhibitors, Monoclonal antibodies (other), Immunoglobulins/Plasma, Antithrombotics, Antimicrobials, Immunomodulators (non–steroid), Kinase inhibitors, Metabolic agents, Supportive care, Control/Standard care, and Other/Unknown. At the bucket–month level we compute two endpoints: (i) Clinical Improvement @ D28 (successes == number improved; trials == reported denominator, or baseline $N$ if missing) and (ii) Survival @ D28 derived from mortality (successes == denominator - deaths). To form a long non–stationary sequence, we adopt a union construction: for each $(k,t,\text{endpoint})$ bin we emit exactly $s_{k,t}$ ones and $n_{k,t}{-}s_{k,t}$ zeros and concatenate all bins in a fixed order (month, clinD28, mortD28, bucket). The sequence is fully deterministic; in our run it comprises $T\!\approx\!7.4{\times}10^{4}$} rounds with $13$ actions.
Figure 7: Raw rewards for COVID-NMA Clinical dataset [10].
Yahoo! R6A Dataset Construction.
For the NS bandit benchmark based on the Yahoo! R6A click log dataset [^8] Yahoo! Front Page Today Module User Click Log Dataset: https://webscope.sandbox.yahoo.com., we follow the main procedure provided in [13, 69]. We merge ten consecutive days of logs and we group the data by article ID and compute smoothed click-through rates (CTRs) using centered rolling averages over a 100-round window. This generates a time series of empirical CTRs for each article. We segment the dataset into ten distinct subperiods (each spanning half a day), filtering out actions with missing data or high noise. We further select a set of common actions present in all segments to ensure consistent tracking. We average CTRs within each subperiod and smoothing small deviations below a threshold $0.005$. We stack selected actions across multiple days into a single $K\times T$ matrix, where $K$ is the number of valid actions and $T$ the compressed time horizon. To reduce spurious noise and compress the time scale, we apply local smoothing. Finally, we apply post-processing filters to remove (i) globally high-value actions (outliers with inflated CTRs), and (ii) actions that persist as best for too many segments.
Figure 8: Mean rewards for the Yahoo! R6A dataset.
Yahoo! R6B Dataset Construction.
We follow a two-stage pipeline tailored to the Yahoo! R6B logs.8 Stage 1 (action vocabulary): we scan the logs to count displays and clicks per article and select the top items using the click-through rate with a minimum display threshold of 2, yielding a fixed action set with mapping $\text{id}\!\mapsto\!k\in\{0,\dots,K-1\}$ with $K=51$, chosen on the same window as the evaluation files. Stage 2 (replay log): we reprocess the files and, for each round $t$, form a feature vector $\mathbf{x}_{t}$ from the given features, restrict the candidate set to the Top–$K$ vocabulary to obtain $\mathcal{A}_{t}$, locate the displayed item’s index $j_{t}^{\star}\in\{0,\dots,|\mathcal{A}_{t}|-1\}$, and record the binary click $X_{t}\in\{0,1\}$; we drop rounds where the displayed item lies outside Top–$K$ or $|\mathcal{A}_{t}|<2$. To increase coverage at a fixed horizon $T=50000$, days are merged in a round-robin fashion before truncation. The resulting dataset stores $\{\mathbf{x}_{t},\mathcal{A}_{t},j_{t}^{\star},r_{t},t_{t}\}_{t=1}^{T}$. For offline replay evaluation, a policy $\pi$ observes $(\mathbf{x}_{t},\mathcal{A}_{t})$ and proposes $A_{t}\in\{0,\dots,|\mathcal{A}_{t}|-1\}$; we credit the outcome only when $a_{t}=j_{t}^{\star}$, and report cumulative reward $C_{T}=\sum_{t=1}^{T}\mathds{1}\{a_{t}=j_{t}^{\star}\}r_{t}$.
Figure 9: Rewards for the Yahoo! R6B dataset.
@@
We use the Bioliq dataset from [34], comprising a week of readings from 20 power plant sensors. Following their setup, we construct an NS-SCB environment with 190 actions: the reward is $1$ if the last 1000 measurements exceed 2.04, and 0 otherwise. Evaluation is based on cumulative reward. Data available at https://github.com/edouardfouche/G-NS-MAB/tree/master/data.
@@
We evaluate on the Zozo environment, a real-world ad recommender system from [51], using the preprocessed dataset of [34]. We construct an NS-GLB environment with all 80 ads (unlike their 10-action setup), assigning reward $1$ to any ad clicked within one second, and 0 otherwise. Evaluation is based on cumulative reward. Data available at https://github.com/edouardfouche/G-NS-MAB/tree/master/data.
@@
Live Traffic Data Construction.
We construct a NS bandit environment based on the Criteo live traffic dataset [20], following the preprocessing approach of [50] but modeling the problem as an NS-GLB rather than an NS-LB. Specifically, the dataset includes banners shown to users, associated contextual variables, and whether each banner was clicked. We retain the categorical variables cat1 through cat9, along with campaign, which uniquely identifies each campaign. These categorical features are one-hot encoded, and a dimensionality reduction via Singular Value Decomposition selects 50 resulting features. The parameter vector $\theta^{\star}$ is estimated using logistic regression. Rewards are then generated from this regression model with added Gaussian noise of variance $\sigma^{2}=0.01$. Unlike [50], in which the authors employ a single change, we introduce shifts in $\theta^{*}$ via a geometric change-point model with parameter $\xi=0.8$, chaging $60\%$ of the $\theta^{*}$ coordinates at each time-step to $\theta^{*}$ and extend the horizon to $T=50000$.